# A kernel conditional two-sample test

### v1

@@ F.2.1 Global and local sensitivity
Noise $\mathcal{N}(0,0.1^{2})$
Bootstrap resamples $500$
@@
Noise $\mathcal{N}(0,0.025^{2})$
Bootstrap resamples $500$
F.2.2 Higher-order moment matching
Figure 3 shows the positive rate of our test for different kernels $\kappa$ on $\mathcal{Z}=\mathbb{R}$ when the conditional means coincide, but the conditional distributions still differ. To that end, we draw two data sets $D_{1}$ and $D_{2}$, each of size $N=100$, by first sampling covariates uniformly from $\mathcal{X}=[-1,1]^{2}$ and evaluating a single shared, randomly chosen unit-norm mean function $f=f_{1}=f_{2}\in\mathcal{H}_{k}$. In $D_{1}$, observations are corrupted by homoscedastic Gaussian noise $\mathcal{N}(0,\sigma^{2})$ with $\sigma=0.025$. In $D_{2}$, we instead add noise drawn from an equally weighted mixture of $\mathcal{N}(-\mu,\sigma^{2})$ and $\mathcal{N}(\mu,\sigma^{2})$. Because both noise models have zero mean, the conditional expectations coincide exactly, yet as $\mu$ increases the variance, skewness, and kurtosis of the mixture diverge from those of the single Gaussian.
We report any additional hyperparameter values in Table 4. The experiment was conducted on an Intel Xeon 8468 Sapphire CPU, using 10 GB of RAM and 20 core hours.
Table 4: Hyperparameters used in generating Figure 3.
Parameter Value
Mean function dimension $m=12$
Regularization $\lambda=0.5$
Bootstrap resamples $500$
F.2.3 Influence of key parameters
We evaluate how different key parameters affect both type $\mathrm{I}$ and type $\mathrm{II}$ error rates by varying one parameter at a time while holding all others fixed. To estimate type $\mathrm{I}$ error, we generate each pair of data sets from the same randomly drawn unit-norm mean function $f\in\mathcal{H}_{k}$, apply i.i.d. Gaussian noise, and record the empirical positive rate. To estimate type $\mathrm{II}$ error, we draw two independent unit-norm mean functions $f_{1},f_{2}\in\mathcal{H}_{k}$, add the same noise model, and treat any rejection at any covariate as a true positive, ignoring the negligible event $f_{1}(x)=f_{2}(x)$. In all cases, covariates are sampled uniformly from $\mathcal{X}=[-1,1]^{2}$. We fix the data set size to $N=100$ and use the inhomogeneous linear kernel on $\mathcal{Z}=\mathbb{R}$ except when we vary those parameters to study their impact.
Output kernel
In Figure 4, we vary the output kernel $\kappa$, comparing different Gaussian and polynomial kernels.
Analytical versus bootstrapped
In Figure 7, we compare the error rates between analytical and bootstrapped thresholds for varying noise levels.
Data set size
Figure 6, we compare the error rates resulting from bootstrapped thresholds for varying data set sizes $N$.
Tables 5, 6, and 7 report any additional hyperparameters for the above experiments. The experiments were conducted on an Intel Xeon 8468 Sapphire CPU, using 10 GB of RAM and 40 core hours.
Table 5: Hyperparameters used in generating Figure 4.
Parameter Value
Mean function dimension $m=12$
Regularization $\lambda=0.5$
Noise $\mathcal{N}(0,0.2^{2})$
Bootstrap resamples $500$
Table 6: Hyperparameters used in generating Figure 4.
Parameter Value
Mean function dimension $m=12$
Regularization $\lambda=0.25$
Bootstrap resamples $500$
Table 7: Hyperparameters used in generating Figure 6.
Parameter Value
Mean function dimension $m=12$
Regularization $\lambda=0.5$
Noise $\mathcal{N}(0,0.2^{2})$
Bootstrap resamples $500$
F.3 Process monitoring
Table 8 presents any additional hyperparameter values not already detailed in Section A.3. The experiment was conducted on an Intel Xeon 8468 Sapphire CPU, using 10 GB of RAM and 2000 core hours.
@@ F.3 Process monitoring
Bootstrap resamples $100$

### v2

@@ G.3 Process monitoring (Figure 3)
Bootstrap resamples $M=500$
@@
Bootstrap resamples $M=500$
@@
Bootstrap resamples $M=100$
Table 6: Hyperparameters used in generating Figure 4.
Parameter Value
Input set $\mathcal{X}=[-1,1]^{2}$
Gaussian output kernel bandwidth $\gamma^{2}=0.05$
Mean function dimension $m=12$
Data set size $n=100$
Noise variance $s^{2}=0.025^{2}$
Regularization $\lambda=0.5$
Bootstrap resamples $M=500$
Table 7: Hyperparameters used in generating Figure 5.
Parameter Value
Input set $\mathcal{X}=[-1,1]^{2}$
Mean function dimension $m=12$
Noise variance $s^{2}=0.2^{2}$
Regularization $\lambda=0.5$
Bootstrap resamples $M=500$
Table 8: Hyperparameters used in generating Figure 6(a) and 6(b).
Parameter Value
Input set $\mathcal{X}=[-1,1]^{2}$
Mean function dimension $m=12$
Noise variance $s^{2}=0.2^{2}$
Bootstrap resamples $M=500$
Table 9: Hyperparameters used in generating Figure 7.
Parameter Value
Input set $\mathcal{X}=[-1,1]^{2}$
Mean function dimension $m=12$
Data set size $n=100$
Regularization $\lambda=0.25$
Bootstrap resamples $M=500$